OpenStudy (derpisthewordorisit):

Select the mathematical sentence that is true. A. 3 • (7 – 3) + 5 > 6 + 2 • 1 B. 3 • (7 – 3) + 5 ≤ 6 + 2 • 1 C. 3 • (7 – 3) + 5 = 6 + 2 • 1 D. 3 • (7 – 3) + 5 < 6 + 2 • 1 @forgiveness

OpenStudy (anonymous):

A is true because: 3 x 4 + 5 = 12 + 5 = 17

OpenStudy (derpisthewordorisit):

Okay, thanks to you I'm starting to get it a little bit.

OpenStudy (anonymous):

17 > 6 + 2 x 1 because: 6 + 2 17 > 8